0.22 of a divalent metal is deposited the current of 0.45A Is passed through a solution of it's salt for 25 minute using appropriate electrodes. Calculate the relative atomic mass of the matal (if=96500C)
1
Expert's answer
2021-09-26T12:15:05-0400

According to the Faraday's laws of electrolysis, the quantity of metal "n" deposited on the electrode is proportional to the quantity of charge "Q = It" which has flown in electrolysis time "t" when a current "I" is passed through the solution:

"n = \frac{It}{zF}" ,

where "z" is the valence of the metal and "F" is the Faraday's constant, equal to 96487 C/mol.

Therefore, the number of the moles of the metal deposited when a current of 0.45 A is passed through a solution of its salt for 25 minutes (25·60 seconds) is:

"n = \frac{0.45 \text{A}\cdot25\cdot60\text{s}}{2·96487\text{C/mol}} = 3.50·10^{-3}" mol.

The relative atomic mass of the metal is the ratio of its mass to its number of the moles:

"M = \frac{m}{n} = \frac{0.222\text{g}}{3.50·10^{-3}\text{mol}} = 63.5" g/mol.

Answer: the relative atomic mass of the metal is 63.5 g/mol. This metal is copper Cu.